### v1

# Spectral Truncation Kernels: Noncommutativity in $C^{*}$-algebraic Kernel Machines

Abstract
In this paper, we propose a new class of positive definite kernels based on the spectral truncation, which has been discussed in the fields of noncommutative geometry and $C^{*}$-algebra. We focus on kernels whose inputs and outputs are functions and generalize existing kernels, such as polynomial, product, and separable kernels, by introducing a truncation parameter $n$ that describes the noncommutativity of the products appearing in the kernels. When $n$ goes to infinity, the proposed kernels tend to the existing commutative kernels. If $n$ is finite, they exhibit different behavior, and the noncommutativity induces interactions along the data function domain. We show that the truncation parameter $n$ is a governing factor leading to performance enhancement: by setting an appropriate $n$, we can balance the representation power and the complexity of the representation space. The flexibility of the proposed class of kernels allows us to go beyond previous commutative kernels.
1 Introduction
Kernel methods have been one of the most fundamental tools in machine learning [1, 2, 3, 4, 5]. They have been applied, for example, to ridge regression, principal component analysis, and support vector machine. Kernel methods are characterized by function spaces called reproducing kernel Hilbert spaces (RKHSs), which are constructed by positive definite kernels. Typical positive definite kernels include the polynomial kernel, Gaussian kernel, and Laplacian kernel. Product kernels, which are constructed by the product of multiple kernels, have also been considered [1, 6].
Standard positive definite kernels are real- or complex-valued, and are well-suited to learn scalar-valued functions. Kernel methods for vector-, function-, or operator-valued outputs have also been investigated [4, 7, 8]. The kernel, in this case, is instead an operator- or function-valued positive definite function, and the associated feature space is a vector-valued RKHS (vvRKHS) [9, 7, 10] or a reproducing kernel Hilbert $C^{*}$-module (RKHM) [8]. RKHM is a generalization of RKHS by means of $C^{*}$-algebra, and it also generalizes vvRKHS. $C^{*}$-algebra is a generalization of the space of complex values and has a structure of the norm, product, and involution [11, 12]. It unifies operators and functions. Thus, the operator- and function-valued kernels for vvRKHSs and RKHMs can be unified as $C^{*}$-algebra-valued kernels. It has been shown that the norms in $C^{*}$-algebras alleviate the dependency of the generalization bound on the output dimension [13] and the product structure of a $C^{*}$-algebra goes beyond the convolution operation [14].
One big challenge for kernel methods is the choice of positive definite kernels. Especially for vvRKHSs and RKHMs, the construction of good $C^{*}$-algebra-valued kernels is not obvious. A typical kernel is the separable kernel, which is defined by the product of a real- or complex-valued kernel and a positive semi-definite operator [4]. Attempts have been made to construct positive definite kernels that go beyond separable kernels. Huusari and Kadri [15] proposed entangled kernels based on concepts from quantum computing, such as partial trace and entanglement. Hashimoto et al. [14] proposed to use the product of circulant matrices and general squared matrices to construct positive definite kernels. By using the kernel, we can generalize the convolution and capture the effect of interactions of different Fourier components.
In this paper, we propose a new class of $C^{*}$-algebra-valued positive definite kernels based on the spectral truncation, which has been discussed in the field of noncommutative geometry and $C^{*}$-algebra [16, 17, 18]. We focus on the case of functional inputs and outputs, and consider the $C^{*}$-algebra of continuous functions. We generalize typical existing kernels: polynomial, product, and separable kernels using the pair of maps called the approximate order isomorphism, which takes an important role in the spectral truncation [19, 17, 20]. The proposed kernels are parameterized by a natural number $n$ that corresponds to the dimension of the truncated space. We approximate functions on the $n$-dimensional truncated space and obtain $n$ by $n$ Toeplitz matrices. Thus, $n$ describes the resolution of the discretization of the function, and we call it the truncation parameter. We show that the proposed kernels converge to the existing kernels as $n$ goes to infinity.
To analyze functional data, we have been focused on how to interpolate observed data to obtain functions [21, 22, 8]. On the other hand, our results shed light on good effects of the discretization (setting $n$ as a finite number) on learning process. Indded, the truncation parameter $n$ takes an important role from at least two perspectives. First, one of the important structures in the existing kernels is the product, and $n$ describes the noncommutativity of the product. Although the product of two functions is too simple to capture structures of these functions, the product of two approximated Toeplitz matrices is more complex, which enables us to capture the structures of the functions. More precisely, for two functions $a$ and $b$, their product $ab$ is defined as just the pointwise product of each value, i.e., $ab(z)=a(z)b(z)$, and the product is commutative, i.e., $ab=ba$. Therefore, we cannot capture the relationship between $a(z_{1})$ and $b(z_{2})$ for $z_{1}\neq z_{2}$ with this product. On the other hand, the product of two Toeplitz matrices is not commutative, and it can be shown that this product involves the product of $a(z_{1})$ and $b(z_{2})$ for $z_{1}\neq z_{2}$. This enables us to capture the relationship between $a(z_{1})$ and $b(z_{2})$. Second, the parameter $n$ controls the tradeoff between the representation power and the model complexity of the RKHM. We show that if $n$ is small, then the representation power is low, and the model complexity is small. On the other hand, if $n$ is large, then the representation power is high, and the model complexity is large. The expectation loss for learning a function is decomposed into the empirical loss, controlled by the representation power, and the complexity. By introducing the parameter $n$ and setting $n$ to balance the empirical loss and the complexity, we obtain higher performance compared to the case of existing kernels.
Our contributions are summarized as follows:
  • We propose a new class of function-valued positive definite kernels that generalize existing kernels: polynomial, product, and separable kernels. The proposed kernels are based on the spectral truncation and are indexed by a truncation parameter $n$, which allows us to introduce noncommutativity into the learning process. The proposed kernels induce the interactions along the data function domain. (Definition 6 and Subsection 3.1)
  • We derive a generalization bound for the learning problem in the RKHMs associated with the proposed kernels. The bound implies that the parameter $n$ controls the tradeoff between the representation power and the model complexity of the RKHMs. (Section 4)
  • We apply the proposed kernels to kernel ridge regression. The result illustrates the advantage of introducing the truncation parameter $n$ in practical settings. (Section 5)
We emphasize that although we focus on the case where the inputs and outputs are functions and three types of kernels, our idea of constructing new positive definite kernels is general. We establish a new framework that allows us to go beyond the usual commutative kernels and to assign a significance to noncommutative structures by discretizing functions.
2 Preliminaries
We review mathematical notions required in this paper.
2.1 $C^{*}$-algebra and Reproducing kernel Hilbert $C^{*}$-module
$C^{*}$-algebra is a Banach space equipped with a product and an involution satisfying the $C^{*}$ identity (condition 3 below). It is a natural generalization of the space of complex numbers.
Definition 1 ($C^{*}$-algebra)
A set $\mathcal{A}$ is called a $C^{*}$-algebra if it satisfies the following conditions:
  1. 1.
    $\mathcal{A}$ is an algebra over $\mathbb{C}$ and equipped with a bijection $(\cdot)^{*}:\mathcal{A}\to\mathcal{A}$ that satisfies the following conditions for $\alpha,\beta\in\mathbb{C}$ and $a,b\in\mathcal{A}$:
    $\bullet$ $(\alpha a+\beta b)^{*}=\overline{\alpha}a^{*}+\overline{\beta}b^{*}$,  $\bullet$ $(ab)^{*}=b^{*}a^{*}$,  $\bullet$ $(a^{*})^{*}=a$.
  2. 2.
    $\mathcal{A}$ is a normed space endowed with $\|\cdot\|_{\mathcal{A}}$, and for $a,b\in\mathcal{A}$, $\|ab\|_{\mathcal{A}}\leq\|a\|_{\mathcal{A}}\|b\|_{\mathcal{A}}$ holds. In addition, $\mathcal{A}$ is complete with respect to $\|\cdot\|_{\mathcal{A}}$.
  3. 3.
    For $a\in\mathcal{A}$, the $C^{*}$ identity $\|a^{*}a\|_{\mathcal{A}}=\|a\|_{\mathcal{A}}^{2}$ holds.
If there exists $a\in\mathcal{A}$ such that $ab=b=ba$ for any $b\in\mathcal{A}$, $a$ is called the unit, and denoted by $1_{\mathcal{A}}$.
In this paper, we focus on the $C^{*}$-algebra of continuous functions.
Example 1
Let $\mathbb{T}=\mathbb{R}/2\pi\mathbb{Z}$ be the torus and $C(\mathbb{T})$ be the space of continuous functions on $\mathbb{T}$. Then, $\mathcal{A}:=C(\mathbb{T})$ is a $C^{*}$-algebra by means of the product: $(cd)(z)=c(z)d(z)$, involution: $c^{*}(z)=\overline{c(z)}$, and norm: $\|c\|_{\mathcal{A}}=\sup_{z\in\mathbb{T}}|c(z)|$ for $c,d\in\mathcal{A}$.
We review basic notions regarding $C^{*}$-algebra. In the following, let $\mathcal{A}$ be a $C^{*}$-algbra.
Definition 2 (Positive)
An element $a$ of $\mathcal{A}$ is called positive if there exists $b\in\mathcal{A}$ such that $a=b^{*}b$ holds. For $a,b\in\mathcal{A}$, we write $a\leq_{\mathcal{A}}b$ if $b-a$ is positive. We denote by $\mathcal{A}_{+}$ the subset of $\mathcal{A}$ composed of all positive elements in $\mathcal{A}$.
Definition 3 (Infimum and minimum)
For a subset $\mathcal{S}$ of $\mathcal{A}$, $a\in\mathcal{A}$ is said to be a lower bound with respect to the order $\leq_{\mathcal{A}}$, if $a\leq_{\mathcal{A}}b$ for any $b\in\mathcal{S}$. Then, a lower bound $c\in\mathcal{A}$ is said to be an infimum of $\mathcal{S}$, if $a\leq_{\mathcal{A}}c$ for any lower bound $a$ of $\mathcal{S}$. If $c\in\mathcal{S}$, then $c$ is said to be a minimum of $\mathcal{S}$.
We now define RKHM. Let $\mathcal{X}$ be a non-empty set for data. To construct an RKHM, we first introduce $\mathcal{A}$-valued positive definite kernel.
Definition 4 ($\mathcal{A}$-valued positive definite kernel)
An $\mathcal{A}$-valued map $k:\mathcal{X}\times\mathcal{X}\to\mathcal{A}$ is called a positive definite kernel if it satisfies the following conditions:
$\bullet$ $k(x,y)=k(y,x)^{*}$  for $x,y\in\mathcal{X}$,
$\bullet$ $\sum_{i,j=1}^{n}c_{i}^{*}k(x_{i},x_{j})c_{j}$ is positive semi-definite  for $n\in\mathbb{N}$, $c_{i}\in\mathcal{A}$, $x_{i}\in\mathcal{X}$.
Let $\phi:\mathcal{X}\to\mathcal{A}^{\mathcal{X}}$ be the feature map associated with $k$, defined as $\phi(x)=k(\cdot,x)$ for $x\in\mathcal{X}$ and let $\mathcal{M}_{k,0}=\{\sum_{i=1}^{n}\phi(x_{i})c_{i}|\ n\in\mathbb{N},\ c_{i}\in\mathcal{A},\ x_{i}\in\mathcal{X}\ (i=1,\ldots,n)\}$. We can define an $\mathcal{A}$-valued map $\big{\langle}\cdot,\cdot\big{\rangle}_{k}:\mathcal{M}_{k,0}\times\mathcal{M}_{k,0}\to\mathcal{A}$ as
$\bigg{\langle}\sum_{i=1}^{n}\phi(x_{i})c_{i},\sum_{j=1}^{l}\phi(y_{j})d_{j}\bigg{\rangle}_{k}=\sum_{i=1}^{n}\sum_{j=1}^{l}c_{i}^{*}k(x_{i},y_{j})d_{j},$
which enjoys the reproducing property $\left\langle\phi(x),f\right\rangle_{k}=f(x)$ for $f\in\mathcal{M}_{k,0}$ and $x\in\mathcal{X}$.
The norm in $\mathcal{M}_{k,0}$ is defined as $\|f\|_{k}^{2}=\|\left\langle f,f\right\rangle_{k}\|_{\mathcal{A}}$, and we can also define the $\mathcal{A}$-valued absolute value $|\cdot|_{k}$ as $|f|_{k}^{2}=\left\langle f,f\right\rangle_{k}$. The reproducing kernel Hilbert $\mathcal{A}$-module (RKHM) $\mathcal{M}_{k}$ associated with $k$ is defined as the completion of $\mathcal{M}_{k,0}$ with respect to $\|\cdot\|_{k}$. See, for example, the references [12, 11, 23, 8] for more details about $C^{*}$-algebra and RKHM.
2.2 Spectral truncation on the torus
The product in $C(\mathbb{T})$ is commutative. However, by approximating the multiplication of a function $x\in C(\mathbb{T})$ by a Toeplitz matrix, we can obtain a noncommutative product structure. Let $e_{j}(z)=\mathrm{e}^{\mathrm{i}jz}$ for $j\in\mathbb{Z}$ and $z\in\mathbb{T}$ be the Fourier function and $M_{x}$ be the multiplication operator defined on $L^{2}(\mathbb{T})$ with respect to $x$. Here, $\mathrm{i}$ is the imaginary unit, and $L^{2}(\mathbb{T})$ is the space of square integrable functions on $\mathbb{T}$. Let $P_{n}$ be the orthogonal projection onto the $n$-dimensional subspace $\operatorname{Span}\{e_{1},\ldots,e_{n}\}$. We approximate $M_{x}$ by $P_{n}M_{x}P_{n}$, i.e., by truncating the spectrum. Then, the $(j,l)$-entry of the representation matrix $R_{n}(x)\in\mathbb{C}^{n\times n}$ of $P_{n}M_{x}P_{n}$ is written as
$R_{n}(x)_{j,l}=\left\langle e_{j},M_{x}e_{l}\right\rangle_{L^{2}(\mathbb{T})}=\int_{\mathbb{T}}x(t)\mathrm{e}^{-\mathrm{i}(j-l)t}\mathrm{d}t.$
Since the $(j,l)$-entry of $R_{n}(x)$ depends only on $j-l$, $R_{n}(x)$ is a Toeplitz matrix [24]. It is characterized only by a vector $r_{n}(x)\in\mathbb{C}^{2n-1}$, where the $(j-l)$th element of $r_{n}(x)$ is $R_{n}(x)_{j,l}$. Note that $R_{n}(x)_{j,l}$ is regarded as the $(j-l)$th Fourier component of $x$, and the vector $r_{n}(x)$ is the coordinate of the function $x$ projected on the space $\operatorname{Span}\{e_{-(n-1)},\ldots,e_{n-1}\}$.
For a matrix $A\in\mathbb{C}^{n\times n}$, let $S_{n}(A)\in C(\mathbb{T})$ be the function defined as $S_{n}(A)(z)=1/n\sum_{j,l=0}^{n-1}A_{j,l}\mathrm{e}^{\mathrm{i}(j-l)z}$, where $A_{j,l}$ is the $(j,l)$-entry of $A$. The map $S_{n}$ takes the representation matrix $R_{n}(x)$ back to a function that approximates the original function $x$. Indeed, we have
$S_{n}(R_{n}(x))(z)=\int_{\mathbb{T}}x(t)\frac{1}{n}\sum_{j,l=0}^{n-1}\mathrm{e}^{-\mathrm{i}(j-l)t}\mathrm{e}^{\mathrm{i}(j-l)z}\mathrm{d}t=\int_{\mathbb{T}}x(t)\frac{1}{n}\sum_{j=0}^{n-1}\sum_{l=-j}^{j}\mathrm{e}^{\mathrm{i}l(z-t)}\mathrm{d}t=x\ast F_{n}(z),$
where $F_{n}(t)=1/n\sum_{j=0}^{n-1}\sum_{l=-j}^{j}\mathrm{e}^{\mathrm{i}lt}$ is the Fejér kernel and $\ast$ represents the convolution. By the following proposition, for each $z\in\mathbb{T}$, $S_{n}(R_{n}(x))(z)$ converges to $x(z)$ as $n\to\infty$ [25, 17].
Proposition 5
For each $z\in\mathbb{T}$, $x\ast F_{n}(z)\to x(z)$ as $n\to\infty$.
The pair of maps $R_{n}$ and $S_{n}$ is called the approximate order isomorphism.
3 $C^{*}$-algebra-valued Positive Definite Kernel with Spectral Truncation
In the following, we set the $C^{*}$-algebra $\mathcal{A}$ as $C(\mathbb{T})$. The following three kinds of $\mathcal{A}$-valued kernels are often considered in existing studies [14, 13, 1, 9, 4].
Example 2
For $x=[x_{1},\ldots,x_{d}],y=[y_{1},\ldots,y_{d}]\in\mathcal{A}^{d}$ and $z\in\mathbb{T}$,
  1. 1.
    polynomial kernel $k^{\operatorname{poly},q}(x,y)(z)=\sum_{i=1}^{d}\alpha_{i}(\overline{x_{i}(z)}y_{i}(z))^{q}$,
  2. 2.
    product kernel $k^{\operatorname{prod},q}(x,y)(z)=\prod_{j=1}^{q}\overline{\tilde{k}_{1,j}(x(z),y(z))}\tilde{k}_{2,j}(x(z),y(z))$,
  3. 3.
    separable kernel $k^{\operatorname{sep},q}(x,y)=\tilde{k}(x,y)\prod_{j=1}^{q}(a_{j}^{*})^{q}a_{j}^{q}$.
Here, $q\in\mathbb{N}$ is the degree of the products in the kernel, $\alpha_{i}\geq 0$ is the parameter of $k^{\operatorname{poly},q}$, and for $i=1,2$ and $j=1,\ldots,q$, $\tilde{k}_{i,j}:\mathbb{C}\times\mathbb{C}\to\mathbb{C}$ and $\tilde{k}:\mathcal{A}^{d}\times\mathcal{A}^{d}\to\mathbb{C}$ are complex-valued continuous positive definite kernels. In addition, $a_{j}\in\mathcal{A}$ is the parameter of $k^{\operatorname{sep},q}$.
We propose new $\mathcal{A}$-valued positive definite kernels by generalizing the existing kernels. The kernels in Example 2 involve the product of functions in $\mathcal{A}$. For $x,y\in\mathcal{A}$, the product of $x$ and $y$ is commutative and is just the pointwise product defined as $(xy)(z)=x(z)y(z)$ (See Example 1). Thus, the kernels defined in Examples 2.1 and 2.2 do not induce the interactions along $z\in\mathbb{T}$, that is, $k(x,y)(z)$ is determined only by the values $x(z)$ and $y(z)$ at $z$. If the values $x(z_{1})$ and $x(z_{2})$, or $x(z_{1})$ and $y(z_{2})$ for different $z_{1}$ and $z_{2}$ are related, then the construction of the kernels only with this commutative product cannot extract that relationship. On the other hand, if we introduce the truncation parameter $n$ and transform $x$ and $y$ into Toeplitz matrices using the map $R_{n}$ defined in Subsection 2.2, then the product of $R_{n}(x)$ and $R_{n}(y)$ is noncommutative. Focusing on this feature, we define $\mathcal{A}$-valued kernels based on Example 2, but they have interactions along $z\in\mathbb{T}$.
Definition 6
With the notations in Example 2, for $x=[x_{1},\ldots,x_{d}],y=[y_{1},\ldots,y_{d}]\in\mathcal{A}^{d}$ and $z\in\mathbb{T}$, let
$k_{n}^{\operatorname{poly},q}(x,y)$ $=S_{n}\bigg{(}\sum_{i=1}^{d}\alpha_{i}(R_{n}(x_{i})^{*})^{q}R_{n}(y_{i})^{q}\bigg{)},$
$k_{n}^{\operatorname{prod},q}(x,y)$ $=S_{n}\bigg{(}\prod_{j=1}^{q}R_{n}(\tilde{k}_{1,j}(x,y))^{*}\prod_{j=1}^{q}R_{n}(\tilde{k}_{2,j}(x,y))\bigg{)},$
$k_{n}^{\operatorname{sep},q}(x,y)$ $=\tilde{k}(S_{n}(R_{n}(x)),S_{n}(R_{n}(y)))S_{n}\bigg{(}\prod_{j=1}^{q}R_{n}(a_{j})^{*}\prod_{j=1}^{q}R_{n}(a_{j})\bigg{)}.$
Here, $A^{*}$ for a matrix $A$ is the adjoint, and we denote by $\tilde{k}_{i,j}(x,y)$ the map $z\mapsto\tilde{k}_{i,j}(x(z),y(z))$.
In Definition 6, to construct the kernels, we first project functions in $\mathcal{A}$ onto the space $\operatorname{Span}\{e_{-(n-1)},\ldots,e_{n-1}\}$ using the map $R_{n}$ and obtain Toeplitz matrices. Then, we consider the product of the Toeplitz matrices. Note that the product of two Toeplitz matrices is not always a Toeplitz matrix. Finally, we apply $S_{n}$ to the matrix and take it back to the space $\operatorname{Span}\{e_{-(n-1)},\ldots,e_{n-1}\}$ and obtain the output function of the kernel. Figure 1 schematically shows the construction of the simplest kernel $k_{n}^{\operatorname{poly},1}(x,y)=S_{n}(R_{n}(x)^{*}R_{n}(y))$ defined in Definition 6.
Figure 1: Overview of the construction of the simplest kernel $k_{n}^{\operatorname{poly},1}(x,y)=S_{n}(R_{n}(x)^{*}R_{n}(y))$
By the definitions of $S_{n}$ and $R_{n}$, we can show the following identity for $x_{1},\ldots,x_{q},y_{1},\ldots,y_{q}\in\mathcal{A}$ (see the proof of Theorem 7 in Appendix A):
$S_{n}\bigg{(}\prod_{j=1}^{q}R_{n}(x_{j})^{*}\prod_{j=1}^{q}R_{n}(y)\bigg{)}(z)=\int_{\mathbb{T}^{2q}}\overline{x_{1}(t_{1})}\cdots\overline{x_{q}(t_{q})}y_{1}(t_{q+1})\cdots y_{q}(t_{2q})F_{n}^{q,P}(z\mathbf{1}-t)\mathrm{d}t,$ (1)
[…]
$F_{n}^{q,P}(t)=1/n\sum_{j=1}^{n-1}\sum_{r\in jP\bigcap\mathbb{Z}^{q}}\mathrm{e}^{\mathrm{i}r\cdot t}$ (2)
for $t\in\mathbb{T}^{q}$ is the Fejér kernel on $\mathbb{T}^{q}$ with respect to the polyhedron $P$. Here, for $j\in\mathbb{Z}$, $jP=\{jz\,\mid\,z\in P\}$. The generalization of the Fejér kernel on $\mathbb{T}$, which is defined in Subsection 2.2, to that on $\mathbb{T}^{q}$ with respect to the sum over a polyhedron has been theoretically investigated [26, 25]. Note that $F_{n}^{q,P}(t)$ is a real-valued function since for $r\in P$, we have $-r\in P$. Thus, if a kernel in Example 2 is real-valued, then the corresponding kernel in Definition 6 is also real-valued. Eq. (1) implies that the value of $S_{n}\big{(}\prod_{j=1}^{q}R_{n}(x_{j})^{*}\prod_{j=1}^{q}R_{n}(y_{j})\big{)}$ at $z$ is determined not only by $x_{1}(z),\ldots,x_{q}(z),y_{1}(z),\ldots,y_{q}(z)$, but also by $x_{1}(t_{1}),\ldots,x_{q}(t_{q}),y_{1}(t_{q+1}),\ldots,y_{q}(t_{2q})$ with different values $t_{1},\ldots,t_{2q}$ from $z$.
3.1 Convergence and interactions
We first show that $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep},q}$ defined in Definition 6 converge to $k^{\operatorname{poly},q}$, $k^{\operatorname{prod},q}$, and $k^{\operatorname{sep},q}$ defined in Example 2, respectively. In this sense, the kernels $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep},q}$ are generalizations of existing kernels. In addition, these convergences imply that the interactions along $z\in\mathbb{T}$ in the kernels become smaller as $n$ grows since they disappear at $n=\infty$. In the following, all the proofs of theoretical statements are documented in Appendix A.
Theorem 7
For $x,y\in\mathcal{A}^{d}$ and $z\in\mathbb{T}$, $k_{n}^{\operatorname{poly},q}(x,y)(z)\to k^{\operatorname{poly},q}(x,y)(z)$, $k_{n}^{\operatorname{prod},q}(x,y)(z)\to k^{\operatorname{prod},q}(x,y)(z)$ as $n\to\infty$. If $x$ and $y$ are differentiable, then $k_{n}^{\operatorname{sep},q}(x,y)(z)\to k^{\operatorname{sep},q}(x,y)(z)$.
As we can see in Eq. (1), if the kernel involves $q$ products of Toeplitz matrices generated by $R_{n}$, it is represented by a Fejér kernel on $\mathbb{T}^{q}$. To show Theorem 7, we apply the following lemmas.
Lemma 8
[…]
Lemma 9
Let $P$ be a convex polyhedron. Let $F_{n}^{q,P}$ be the Fejér kernel on $\mathbb{T}^{q}$ defined as Eq. (2). Then, for any $g\in C(\mathbb{T}^{q})$ and any $z\in\mathbb{T}^{q}$, $g\ast F_{n}^{q,P}(z)\to g(z)$ as $n\to\infty$.
Since the product of Toeplitz matrices in $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep},q}$ are represented by the sum over the indices in $\bigcup_{j=0}^{2q}Q_{j}^{m}$, Lemma 8 explains why the Fejér kernel with the polyhedron $P=\{r\in\mathbb{R}^{2q}\,\mid\,|\sum_{i=l}^{k}r_{i}|\leq 1,\ l\leq k\}$ appears in Eq. (1). Lemma 9 generalizes the convergence of the Fejér kernel on $\mathbb{T}$ to that on $\mathbb{T}^{q}$, which is by Brandolini and Travaglini [25]. We set $g(z)=\overline{x_{1}(z_{1})}\cdots\overline{x_{q}(z_{q})}y_{1}(z_{q+1})\cdots y_{q}(z_{2q})$ for $z=[z_{1},\ldots,z_{2q}]$, based on Eq. (1), and apply Lemmas 8 and 9 to show Theorem 7. See Appendix A for more details.
As seen in Theorem 7, the interactions along $z\in\mathbb{T}$ in the kernels become small as $n$ grows. On the other hand, we also have the parameter $q$ that describes how many Toeplitz matrices created by $R_{n}$ are involved in the kernel. Regarding the relationship between $q$ and the interactions, we have the following remark.
Remark 10
We expect that the convergence in Theorem 7 becomes slower as $q$ becomes larger, and the interaction along $z\in\mathbb{T}$ becomes larger. This is based on the observation that the value related to the Fejér kernel on $\mathbb{T}^{q}$ is described by the corresponding value related to the Fejér kernel on $\mathbb{T}^{q-1}$, especially in the case of $q=2$. See Appendix B for more details.
3.2 Positive definiteness
To define an RKHM introduced in Subsection 2.1, the kernel $k$ has to be positive definite. Thus, we investigate the positive definiteness of the proposed kernels. Regarding $k_{n}^{\operatorname{poly},q}(x,y)$, we can show the positive definiteness since $x_{i}$ and $y_{i}$ are separated by the product. Regarding $k_{n}^{\operatorname{sep}}(x,y)$, since $x$ and $y$ depend only on $\tilde{k}$, we can use the positive definiteness of $\tilde{k}$.
Proposition 11
The kernels $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{sep}}$ are positive definite.
As for $k_{n}^{\operatorname{prod},q}(x,y)$, $x$ and $y$ depend on $\tilde{k}_{i,j}$, and we cannot separate $x$ and $y$ as products. Thus, we cannot show its positive definiteness. However, we can modify it to become positive definite as follows.
Proposition 12
Let $\beta_{n}\geq-\min_{t\in\mathbb{T}^{q}}F_{n}^{q,P}(t)$. Let
$\hat{k}_{n}^{\operatorname{prod},q}(x,y)=k_{n}^{\operatorname{prod},q}(x,y)+\beta_{n}\int_{\mathbb{T}^{2q}}\prod_{j=1}^{q}\overline{\tilde{k}_{1,j}(x(t_{j}),y(t_{j}))}\tilde{k}_{2,j}(x(t_{q+j}),y(t_{q+j}))\mathrm{d}t.$
Then, $\hat{k}_{n}^{\operatorname{prod},q}$ is positive definite.
To set the value of the parameter $\beta_{n}$, we have the following bound.
Lemma 13
We have $|F_{n}^{q,P}(z)|\leq n^{q}$.
Remark 14
We can set $\beta_{n}$ in $\hat{k}_{n}^{\operatorname{prod},q}$ as $n^{q}$ to guarantee the positive definiteness. However, the bound in Lemma 13 may not be tight, and even if we set a smaller $\beta_{n}$, $\hat{k}_{n}^{\operatorname{prod},q}$ may become positive definite. In addition, considering non-positive kernels has also been investigated [27, 28]. Indeed, in practical computations in Subsection 5.1, the algorithms can work even if we set $\beta_{n}<n^{q}$. Deriving a tighter bound of $\beta_{n}$ or developing the theory for non-positive kernels for RKHMs is future work.
4 Generalization Bound
We now focus on what impact the truncation parameter $n$ has on generalization. We provide a generalization bound allowing us to observe the tradeoff between the representation power and the model complexity of the RKHMs associated with the proposed kernels. To derive the generalization bound, we derive an upper bound of the Rademacher complexity. Let $\Omega$ be a probability space with the probability measure $\mu$. Let $X_{1},\ldots,X_{N}$ and $Y_{1},\ldots,Y_{N}$ be input and output samples from a distributions of $\mathcal{A}_{0}^{d}$-valued and $\mathcal{A}_{1}$-valued random variables $X$ and $Y$ on $\Omega$, respectively (i.e., for $z\in\mathbb{T}$, $X_{i}(z)$ is a sample from the distribution of $X(z)$). Here, $\mathcal{A}_{0}$ and $\mathcal{A}_{1}$ are subsets of $\mathcal{A}$. Let $\sigma_{1},\ldots,\sigma_{N}$ be i.i.d. Rademacher variables, which take their values on $\{-1,1\}$. For a real-valued function class $\mathcal{F}$ and $\mathbf{x}=[x_{1},\ldots,x_{N}]\in(\mathcal{A}^{d})^{N}$, the empirical Rademacher complexity $\hat{R}_{N}(\mathbf{x},\mathcal{F})$ is defined by $\hat{R}_{N}(\mathbf{x},\mathcal{F})=\mathrm{E}[\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}f(x_{i})\sigma_{i}]/N$. Here, $\mathrm{E}$ is the integration on $\Omega$ with respect to $\mu$. An $\mathcal{A}$-valued version of the Rademacher complexity is also defined by Hashimoto et al. [14] as $\hat{R}_{\mathcal{A},N}(\mathbf{x},\mathcal{F})=\mathrm{E}[\sup^{\mathcal{A}}_{f\in\mathcal{F}}\sum_{i=1}^{N}|f(x_{i})^{*}\sigma_{i}|_{\mathcal{A}}]/N$ for an $\mathcal{A}$-valued function class $\mathcal{F}$. Here, $|a|_{\mathcal{A}}=(a^{*}a)^{1/2}$ for $a\in\mathcal{A}$, $\sup^{\mathcal{A}}$ is the supremum in the sense of the order in $\mathcal{A}$ (see Definition 3), and the integral $\mathrm{E}$ means the Bochner integral in this case.
Let $B>0$ and =$\mathcal{F}=\{f\in\mathcal{M}_{k}\,\mid\,\|f\|_{k}\leq B\}$. Let $g:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{+}$ be an error function. Assume there exists $L>0$ such that for $y\in\mathcal{A}_{1}$ and $z\in\mathbb{T}$, $x\mapsto g(x,y(z))$ is $L$-Lipschitz continuous. We derive the following generalization bound for the kernels defined in Section 3.
Theorem 15
Assume $k^{\operatorname{poly},q}$, $k^{\operatorname{prod},q}$, and $k^{\operatorname{sep},q}$ are real-valued. Let $D(k_{n}^{\operatorname{poly},q},x)=\sum_{j=1}^{d}\alpha_{j}\|R_{n}(x_{j})\|^{2q}_{\operatorname{op}}$, $D(\hat{k}_{n}^{\operatorname{prod},q},x)=\prod_{j=1}^{q}(\|R_{n}(\tilde{k}_{1,j}(x,x))\|_{\operatorname{op}}\|R_{n}(\tilde{k}_{2,j}(x,x))\|_{\operatorname{op}})+\beta_{n}C$, and $D(k_{n}^{\operatorname{sep},q},x)=\tilde{k}(x,x)\prod_{j=1}^{q}\|R_{n}(a_{j})\|^{2}$ for $x\in\mathcal{A}_{0}^{d}$, where $\|\cdot\|_{\operatorname{op}}$ is the operator norm and $C=\prod_{j=1}^{q}\int_{\mathbb{T}}\tilde{k}_{1,j}(x(t),x(t))\mathrm{d}t\int_{\mathbb{T}}\tilde{k}_{2,j}(x(t),x(t))\mathrm{d}t$. Assume $\beta_{n}\leq\beta_{n+1}$ for $\hat{k}_{n}^{\operatorname{prod},q}$. For $k_{n}=k_{n}^{\operatorname{poly},q},\hat{k}_{n}^{\operatorname{prod},q},k_{n}^{\operatorname{sep},q}$ and for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
$\mathrm{E}[g(f(X),Y)]\leq_{\mathcal{A}}\frac{1}{N}\sum_{i=1}^{N}g(f(X_{i}),Y_{i})+2L\frac{B}{N}\bigg{(}\sum_{i=1}^{N}D(k_{n},X_{i})\bigg{)}^{1/2}1_{\mathcal{A}}+3\sqrt{\frac{\log 1/\delta}{N}}1_{\mathcal{A}}$ (3)
for any $f\in\mathcal{F}$. In addition, we have $D(k_{n},x)\leq D(k_{n+1},x)$.
The first term of the bound (3) is the empirical loss, and the second term of the bound represents the model complexity of the RKHM $\mathcal{M}_{k_{n}}$ associated with $k_{n}$. There is a tradeoff between these terms. If $n$ is larger, then since the approximation space $\operatorname{Span}\{e_{1},\ldots,e_{n}\}$ in $\mathcal{A}$ is larger, $\mathcal{M}_{k_{n}}$ can represent more functions. Thus, the empirical loss (the first term) can become small if $n$ is large. On the other hand, according to Theorem 15, the complexity of $\mathcal{M}_{k_{n}}$ (the second term) is larger if $n$ is larger. An advantage of introducing the parameter $n$ and constructing the new kernels based on the existing kernels in Example 2 is that we can control the empirical loss and the complexity through $n$. By setting a proper $n$, we can balance the two terms and make the expected loss $\mathrm{E}[g(f(X),Y)]$ small.
To show Theorem 15, we use the $\mathcal{A}$-valued Rademacher complexity. Indeed, in our case, the $\mathcal{A}$-valued Rademacher complexity is represented by the standard Rademacher complexity, and the argument is reduced to evaluate the standard Rademacher complexity.
Proposition 16
Let $\mathcal{F}$ be an $\mathcal{A}$-valued function class. Assume for any $x\in\mathcal{A}_{0}^{d}$ and $z\in\mathbb{T}$, $f(x)(z)\in\mathbb{R}$. Then, we have $\hat{R}_{N}(\mathbf{x},\mathcal{F}(z))=\hat{R}_{\mathcal{A},N}(\mathbf{x},\mathcal{F})(z)$, where $\mathcal{F}(z)=\{x\mapsto(f(x))(z)\,\mid\,f\in\mathcal{F}\}$.
5 Application to Kernel Ridge Regression
We illustrate the effect of the proposed kernels by applying them to kernel ridge regression. Let $x_{1},\ldots,x_{N}\in\mathcal{A}^{d}$ be input training samples and $y_{1},\ldots,y_{N}\in\mathcal{A}$ be output training samples. We consider the RKHM $\mathcal{M}_{k}$ associated with $k=k_{n}^{\operatorname{poly},q}$, $k=k_{n}^{\operatorname{prod},q}$, or $k=k_{n}^{\operatorname{sep},q}$. Consider the typical minimization problem for regression, which is also considered in [14]:
$\min_{f\in\mathcal{M}_{k}}\bigg{(}\sum_{i=1}^{N}|f(x_{i})-y_{i}|_{\mathcal{A}}^{2}+\lambda|f|_{k}^{2}\bigg{)},$ (4)
where $\lambda\geq 0$ is the regularization parameter. We apply the approximation version of the representer theorem for RKHMs over general $C^{*}$-algebras (see Appendix C). Then, the solution of the problem (4) is approximated by a vector having the form of $\sum_{i=1}^{N}\phi_{i}(x_{i})c_{i}$. The problem is reduced to
$\min_{c_{j}\in\mathcal{A}}\bigg{(}\sum_{i=1}^{N}\bigg{|}\sum_{j=1}^{N}k(x_{i},x_{j})c_{j}-y_{i}\bigg{|}_{{\mathcal{A}}}^{2}+\lambda\bigg{|}\sum_{j=1}^{N}\phi(x_{j})c_{j}\bigg{|}_{k}^{2}\bigg{)}$
$\quad=\min_{c_{j}\in\mathcal{A}}(\mathbf{c}^{*}\mathbf{G}^{2}\mathbf{c}-\mathbf{c}^{*}\mathbf{G}\mathbf{y}-\mathbf{y}^{*}\mathbf{G}\mathbf{c}+\lambda\mathbf{c}^{*}\mathbf{G}\mathbf{c}),$ (5)
where $\mathbf{G}$ is the ${\mathcal{A}^{N\times N}}$-valued Gram matrix whose $(i,j)$-entry is defined as $k(x_{i},x_{j}){\in\mathcal{A}}$, $\mathbf{c}=[c_{1},\ldots,c_{N}]^{T}\in\mathcal{A}^{N}$, $\mathbf{y}=[y_{1},\ldots,y_{N}]^{T}\in\mathcal{A}^{N}$. Then, the solution $\mathbf{c}$ of the problem (5) satisfies $\mathbf{y}=(\mathbf{G}+\lambda I)\mathbf{c}$, which means $\mathbf{y}(z)=(\mathbf{G}(z)+\lambda I)\mathbf{c}(z)$. Therefore, for each $z\in\mathbb{T}$, we obtain $\mathbf{c}(z)$ by computing $(\mathbf{G}(z)+\lambda I)^{-1}\mathbf{y}(z)$.
Computational cost
The computational cost for the product of two $n$ by $n$ Toeplitz matrices is $O(n^{2})$, and computing $S_{n}(A)(z)$ for an $n$ by $n$ matrix $A$ and $z\in\mathbb{T}$ costs $O(n^{2})$. Thus, the cost for constructing $\mathbf{G}(z_{1}),\ldots,\mathbf{G}(z_{m})$ for different $m$ points $z_{1},\ldots,z_{m}\in\mathbb{T}$ with $k_{n}^{\operatorname{poly},q}$, $\hat{k}_{n}^{\operatorname{prod},q}$, or $k_{n}^{\operatorname{sep},q}$ is $O((qn^{2}+mn^{2})N^{2})$. As for the computation of $\mathbf{c}(z)$, we need to compute $(\mathbf{G}(z)+\lambda I)^{-1}\mathbf{y}(z)$, which costs $O(N^{3})$. Thus, the total computational cost for obtaining $\mathbf{c}(z_{1}),\ldots,\mathbf{c}(z_{m})$ is $O((q+m)n^{2}N^{2}+mN^{3})$. On the other hand, that for $k^{\operatorname{poly},q}$, $\hat{k}^{\operatorname{prod},q}$, and $k^{\operatorname{sep},q}$ is $O(qmN^{2}+mN^{3})$. Note that the cost is linear with respect to $m$. This is by virtue of considering function-valued kernels and RKHMs. The situation is different if we use a vvRKHS. For vvRKHSs, to obtain $m$ values as an output, we generally apply a $\mathbb{C}^{m\times m}$-valued kernel. In this case, the Gram matrix should be $mN$ by $mN$, which results in a computational cost of $O(m^{3}N^{3})$. We can reduce the cost with respect to $N$, e.g., by Nyström method [29]. Regarding $n$ and $q$, one approach to reducing the cost with respect to them is investigating the method to approximate the Fejér kernel. However, the main contribution of this paper is to show the advantage of introducing the noncommutivity in kernels, and investigating computationally effective methods is future work.
5.1 Numerical results
We show numerical results that illustrate the effects of the noncommutativity on the learning process.
5.1.1 Experiment with synthetic data
To observe the behavior of the proposed kernels $k_{n}^{\operatorname{poly},q}$, $\hat{k}_{n}^{\operatorname{prod},q}$, and ${k}_{n}^{\operatorname{sep},q}$ with different values of $n$ and compare them to the existing kernels $k^{\operatorname{poly},q}$, $k^{\operatorname{prod},q}$, and $k^{\operatorname{sep},q}$ (corresponding to $n=\infty$), we consider a regression task with synthetic data. For $i=1,\ldots,N$, we generated input training samples as $x^{i}(z)=[\sin(0.01iz)+0.01\xi_{z,i},\cos(0.01iz)+0.01\eta_{z,i}]$. We set the target function $f$ as $f(x)(z)=3(\sin(\cos(\int_{z-\Delta}^{z+\Delta}x_{1}(t)\mathrm{d}t+\int_{z-\Delta}^{z+\Delta}x_{2}(t)\mathrm{d}t)))$, where $\Delta=2\pi/30$, and generated training output samples $y^{i}(z)=f(x)(z)+0.001\xi_{z,i}$. Here, $\xi_{z,i}$ and $\eta_{z,i}$ are sampled independently from the Gaussian distribution with mean 0 and standard deviation 1. We set the sample size $N$ as 1000. For the kernel ${k}_{n}^{\operatorname{poly},q}$, we set $q=1$ and $\alpha_{1}=1$. Note that ${k}^{\operatorname{poly},q}$ is the linear kernel in this case. For the kernel $\hat{k}_{n}^{\operatorname{prod},q}$, we set $q=1$, $\tilde{k}_{1,1}(x,y)=\tilde{k}_{2,1}(x,y)=\mathrm{e}^{-|x-y|^{2}}$. We set $\beta_{n}=1$ for $n<\infty$ and $\beta_{\infty}=0$. For the kernel $k_{n}^{\operatorname{sep},q}$, we set $\tilde{k}(x,y)=\mathrm{e}^{-0.1/\Delta^{2}(\|x_{1}-y_{1}\|_{L^{2}(\mathbb{T})}^{2}+\|x_{2}-y_{2}\|_{L^{2}(\mathbb{T})}^{2})}$, $q=2$, and $a_{1}(z)=a_{2}(z)=\mathrm{e}^{\sin z}$. We estimated $f$ using kernel ridge regression. Figure 2 (a-c) illustrates the test error. We can see that for ${k}_{n}^{\operatorname{poly},q}$, the test error is the smallest when $n=32$. It becomes large if $n$ is larger or smaller than $32$. This is because when $n=32$, the empirical error and the complexity in Eq. 3 are balanced, and the best possible expected error is obtained. We can see that even for the simplest kernel ${k}_{n}^{\operatorname{poly},1}$, the proposed kernel ($n$ is finite) goes beyond the existing kernel ($n=\infty$). We have similar results for the other two kernels. Note that $\hat{k}_{\infty}^{\operatorname{prod},q}(x,y)(z)=\hat{k}^{\operatorname{prod},q}(x,y)(z)$ only depends on $x(z)$ and $y(z)$. However, $f(x)(z)$ depends on $x(t)$ for $t\in[z-\Delta,z+\Delta]$. Thus, the test error becomes large when $n=\infty$. In addition, as discussed in Remark 14, $\hat{k}^{\operatorname{prod},q}$ may not be positive definite. However, although we set $\beta_{n}$ as a small value, all the eigenvalues of the Gram matrix are positive in this case. See Appendix D.1 for more details.
(a) $k={k}_{n}^{\operatorname{poly},q}$
(b) $k=\hat{k}_{n}^{\operatorname{prod},q}$
(c) $k=k_{n}^{\operatorname{sep},q}$
(d) $k=\hat{k}_{n}^{\operatorname{prod},q}$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$n=30$ $n=40$ $n=50$ $n=60$ $n=70$ $n=80$ $n=\infty$
(e) Output images
Figure 2: (a-c) Test error of the regression task with different values of $n$. (Box plot of results of five independent runs with different values of noises $\xi_{z,i}$ and $\eta_{z,i}$ in $x^{i}(z)$ and $y^{i}(z)$.) (d) Test error of the image recovering task with different values of $n$. (Box plot of results of five independent runs with different training and test data.) (e) Output images of the image recovering task with different values of $n$.
5.1.2 Experiment with MNIST
We consider an image recovering task with MNIST [30]. For $i=1,\ldots,N$, we generated input training samples by setting all the $8\times 8$ pixels in the middle of the images as 00. We set the output samples as the original images. We flatten the 2-dimensional $28\times 28$ image to a vector in $\mathbb{R}^{784}$ and regard it as a discretized function on the space of pixels. We try to estimate a function that transforms an image with the missing part into the original image. We applied kernel ridge regression with the regularization parameter $\lambda=0.01$. We set the sample size $N$ as $200$, and used the kernel $k_{n}^{\operatorname{prod},q}$ with $q=1$ and $\tilde{k}_{1,1}(x,y)=\tilde{k}_{2,1}(x,y)=\mathrm{e}^{-0.1|x-y|^{2}}$. We set $\beta_{n}=0.01$ for $n<\infty$ and $\beta_{\infty}=0$. Figure 2 (d) and (e) show the test error and the output images with different values of $n$. We can see that the test error becomes the smallest when $n=70$, but it becomes large when $n$ is smaller or larger than $70$. As for the output images, when $n=\infty$, i.e., the existing kernel $k^{\operatorname{prod},q}$, we cannot recover the missing part since each pixel of the output is determined only with the corresponding pixel of the input, and we cannot obtain information of other pixels to estimate what is written in the image. When $n=70$, we can recover the image the clearest.
6 Generalization to Other $C^{*}$-algebras
Setting approximate order isomorphisms for more general $C^{*}$-algebras has been investigated. Using these results and replacing $R_{n}$ and $S_{n}$ in Definition 6, we can define positive definite kernels for the $\mathcal{A}^{d}$-valued inputs and $\mathcal{A}$-valued output for a more general $C^{*}$-algebra $\mathcal{A}$.
Continuous functions on high-dimensional torus [20]
[…]
Continuous functions on the sphere [19]
For $\mathcal{A}=C(S^{2})$, let $\rho_{n}$ be the $n$-dimensional irreducible group representation of $SU(2)$ (special unitary group of degree 2) and let $V_{n}$ be the representation space. Let $P\in\mathcal{B}(V_{n})$ be a projection, where $\mathcal{B}(V_{n})$ is the space of bounded linear operators on $V_{n}$. We set $R_{n}(x)=n\int_{SU(2)}x(g)\rho_{n}(g)^{*}P\rho_{n}(g)\mathrm{d}g$ and $S_{n}(A)(g)=\operatorname{tr}(A\rho_{n}(g)^{*}P\rho_{n}(g))$.
7 Conclusion and Limitation
In this paper, we proposed a new class of positive definite kernels for functional inputs and outputs based on the spectral truncation. The proposed kernels are indexed by a truncation parameter $n$, and they converge to existing kernels: polynomial, product, and separable kernels, as $n$ goes to infinity. Setting $n$ as a finite value, we can introduce noncommutativity into the learning process. More precisely, we can induce interactions along the data function domain and control the tradeoff between the representation power and the model complexity of the RKHMs associated with the proposed kernels. Since $n$ describes the resolution of the discretization functions, our results show the positive effect of discretization on generalization. We applied the proposed kernel to kernel ridge regression and numerically observed that the proposed kernels outperformed the existing kernels.
In the current setting, we focus on the $C^{*}$-algebra of continuous functions on the torus. Although we can generalize this setting and define positive definite kernels for more general inputs and outputs as we discussed in Section 6, theoretical investigations for these generalized settings and their computationally efficient implementations remain to be investigated in future work.
Acknowledgements
The representer theorem stated in Appendix C is pointed out by Ms. Miho Mukohara. We would like to thank her for the helpful discussion. We would also like to thank Mr. Noboru Isobe for a constructive discussion. Hachem Kadri is partially supported by grant ANR-19-CE23-0011 from the French National Research Agency. Masahiro Ikeda is partially supported by grant JPMJCR1913 from JST CREST.
Appendix
Appendix A Proofs
We provide the proofs of theorems, propositions, and lemmas stated in the main text.
Theorem 7    For $x,y\in\mathcal{A}^{d}$ and $z\in\mathbb{T}$, $k_{n}^{\operatorname{poly},q}(x,y)(z)\to k^{\operatorname{poly},q}(x,y)(z)$, $k_{n}^{\operatorname{prod},q}(x,y)(z)\to k^{\operatorname{prod},q}(x,y)(z)$ as $n\to\infty$. If $x$ and $y$ are differentiable, then $k_{n}^{\operatorname{sep},q}(x,y)(z)\to k^{\operatorname{sep},q}(x,y)(z)$.
Proof  We show $k_{n}^{\operatorname{poly},q}(x,y)\to k^{\operatorname{poly},q}(x,y)$. The proofs for $k_{n}^{\operatorname{prod},q}(x,y)$ and $k_{n}^{\operatorname{sep},q}(x,y)$ are the same. We have
$k_{n}^{\operatorname{poly},q}(x,y)=S_{n}\bigg{(}\sum_{i=1}^{d}\alpha_{i}(R_{n}(x_{i})^{*})^{q}R_{n}(y_{i})^{q}\bigg{)}$
$=\frac{1}{n}\sum_{i=1}^{d}\alpha_{i}\sum_{j,l=0}^{n-1}\sum_{r_{1},\ldots,r_{2q}=0}^{n-1}R_{n}(x_{i})^{*}_{j,r_{1}}\cdots R_{n}(x_{i})^{*}_{r_{q-1},r_{q}}R_{n}(y_{i})_{r_{q},r_{q+1}}\cdots R_{n}(y_{i})_{r_{2q-1},l}\mathrm{e}^{\mathrm{i}(j-l)z}$
$=\frac{1}{n}\sum_{i=1}^{d}\alpha_{i}\sum_{j,l=0}^{n-1}\sum_{r_{1},\ldots,r_{2q-1}=0}^{n-1}\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}\overline{x_{i}(t_{1})}\cdots\overline{x_{i}(t_{q})}y_{i}(t_{q+1})\cdots y_{i}(t_{2q})$
$\qquad\cdot\mathrm{e}^{\mathrm{i}(r_{1}-j)t_{1}}\cdots\mathrm{e}^{\mathrm{i}(r_{q}-r_{q-1})t_{q}}\mathrm{e}^{\mathrm{i}(r_{q+1}-r_{q})t_{q+1}}\cdots\mathrm{e}^{\mathrm{i}(l-r_{2q-1})t_{2q}}\mathrm{d}t_{1}\cdots\mathrm{d}t_{2q}\mathrm{e}^{\mathrm{i}(j-l)z}$
$=\frac{1}{n}\sum_{i=1}^{d}\alpha_{i}\sum_{j,l=0}^{n-1}\sum_{r_{1},\ldots,r_{2q-1}=0}^{n-1}\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}\overline{x_{i}(t_{1})}\cdots\overline{x_{i}(t_{q})}y_{i}(t_{q+1})\cdots y_{i}(t_{2q})$
$\qquad\cdot\mathrm{e}^{\mathrm{i}(r_{1}-j)(t_{1}-z)}\cdots\mathrm{e}^{\mathrm{i}(r_{q}-r_{q-1})(t_{q}-z)}\mathrm{e}^{\mathrm{i}(r_{q+1}-r_{q})(t_{q+1}-z)}\cdots\mathrm{e}^{\mathrm{i}(l-r_{2q-1})(t_{2q}-z)}\mathrm{d}t_{1}\cdots\mathrm{d}t_{2q}$
$=\frac{1}{n}\sum_{i=1}^{d}\alpha_{i}\sum_{m=0}^{n-1}\bigg{(}\sum_{j,l,r_{1},\ldots,r_{2q-1}=0}^{m}-\sum_{j,l,r_{1},\ldots,r_{2q-1}=0}^{m-1}\bigg{)}\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}\overline{x_{i}(t_{1})}\cdots\overline{x_{i}(t_{q})}y_{i}(t_{q+1})\cdots y_{i}(t_{2q})$
$\qquad\cdot\mathrm{e}^{\mathrm{i}(r_{1}-j)(t_{1}-z)}\cdots\mathrm{e}^{\mathrm{i}(r_{q}-r_{q-1})(t_{q}-z)}\mathrm{e}^{\mathrm{i}(r_{q+1}-r_{q})(t_{q+1}-z)}\cdots\mathrm{e}^{\mathrm{i}(l-r_{2q-1})(t_{2q}-z)}\mathrm{d}t_{1}\cdots\mathrm{d}t_{2q}$
$=\frac{1}{n}\sum_{i=1}^{d}\alpha_{i}\sum_{m=0}^{n-1}\sum_{\begin{subarray}{c}j\lor l\lor r_{1}\lor\ldots\lor r_{2q-1}=m\\ 0\leq j,l,r_{1},\ldots,r_{2q-1}\leq m\end{subarray}}\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}\overline{x_{i}(t_{1})}\cdots\overline{x_{i}(t_{q})}y_{i}(t_{q+1})\cdots y_{i}(t_{2q})$
$\qquad\cdot\mathrm{e}^{\mathrm{i}(r_{1}-j)(t_{1}-z)}\cdots\mathrm{e}^{\mathrm{i}(r_{q}-r_{q-1})(t_{q}-z)}\mathrm{e}^{\mathrm{i}(r_{q+1}-r_{q})(t_{q+1}-z)}\cdots\mathrm{e}^{\mathrm{i}(l-r_{2q-1})(t_{2q}-z)}\mathrm{d}t_{1}\cdots\mathrm{d}t_{2q}$
$=\frac{1}{n}\sum_{i=1}^{d}\alpha_{i}\sum_{j=0}^{n-1}\sum_{r\in jP\bigcap\mathbb{Z}^{q}}\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}\overline{x_{i}(t_{1})}\cdots\overline{x_{i}(t_{q})}y_{i}(t_{q+1})\cdots y_{i}(t_{2q})$
$\qquad\cdot\mathrm{e}^{\mathrm{i}r_{1}(z-t_{1})}\cdots\mathrm{e}^{\mathrm{i}r_{q}(z-t_{q})}\mathrm{e}^{\mathrm{i}r_{q+1}(z-t_{q+1})}\cdots\mathrm{e}^{\mathrm{i}r_{2q}(z-t_{2q})}\mathrm{d}t_{1}\cdots\mathrm{d}t_{2q}$
$=\sum_{i=1}^{d}\alpha_{i}g_{i}\ast F_{n}^{2q,P}(z),$
where the sum $\sum_{j,l,r_{1},\ldots,r_{2q-1}=0}^{-1}$ is set as 00, $g_{i}(t)=\overline{x_{i}(t_{1})}\cdots\overline{x_{i}(t_{q})}y_{i}(t_{q+1})\cdots y_{i}(t_{2q})$ for $t=[t_{1},\ldots,t_{2q}]\in\mathbb{T}^{2q}$, and $P=\{r=[r_{1},\ldots,r_{2q}]\in\mathbb{R}^{2q}\,\mid\,|\sum_{i=l}^{m}r_{i}|\leq 1,\ l\leq m\}$. In addition, we set $j=r_{0}$ and $j=r_{q}$. The second to the last equality is derived by Lemma 8. Since $P$ is a convex polyhedron, the last equality is by Lemma 9. For $k_{n}^{\operatorname{sep},q}$, we additionally use the fact $\lim_{n\to\infty}\tilde{k}(S_{n}(R_{n}(x)),S_{n}(R_{n}(y)))=\tilde{k}(\lim_{n\to\infty}S_{n}(R_{n}(x)),\lim_{n\to\infty}S_{n}(R_{n}(y)))$, which follows from the continuity of $\tilde{k}$, and the fact $S_{n}(R_{n}(x))\to x$ uniformly as $n\to\infty$ if $x$ and $y$ are differentiable [17, Lemma 10]. $\Box$
[…]
Proof  Let $r^{\prime}\in Q_{j}^{m}$. For $i>j$, we have $r_{i}=r_{i}^{\prime}+r_{i-1}=r_{i}^{\prime}+(r_{i-1}^{\prime}+r_{i-2})=\cdots=m+r_{j+1}^{\prime}+\cdots+r_{i}^{\prime}$, and for $i<j$, we have $-r_{i}=r_{i+1}^{\prime}-r_{i+1}=r_{i+1}^{\prime}+(r_{i+2}^{\prime}-r_{i+2})=\cdots=r_{i+1}^{\prime}+\cdots+r_{j}^{\prime}-m$. Therefore, we have
$\bigcup_{j=0}^{2q}Q_{j}^{m}$
$=\bigcup_{j=0}^{2q}\bigg{\{}r\in\mathbb{R}^{2q}\,\bigg{|}\,r_{i}=\sum_{l=j+1}^{i}r_{l}^{\prime}+m\ (j<i\leq 2q),\ r_{i}=-\sum_{l=i+1}^{j}r_{l}^{\prime}+m\ (0\leq i<j),$
$\qquad\qquad 0\leq r_{i}\leq m\ (i=0,\ldots,j-1,j+1,\ldots 2q)\}$
[…]
=$=\bigg{\{}r\in\mathbb{R}^{2q}\,\bigg{|}\,\bigg{|}\sum_{i=l}^{k}r_{i}\bigg{|}\leq 1,\ l\leq k\bigg{\}},$
which completes the proof of the lemma. $\Box$
Proposition 11    The kernels $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{sep}}$ are positive definite.
Proof  Let $x_{1},\ldots,x_{N}\in\mathcal{A}^{d}$ and $d_{1},\ldots,d_{N}\in\mathcal{A}$. Then, we have
$\bigg{(}\sum_{j,l=1}^{N}d_{j}^{*}k_{n}^{\operatorname{poly},q}(x_{j},x_{l})d_{l}\bigg{)}(z)$
$=\sum_{j,l=1}^{N}\overline{d_{j}(z)}\sum_{j^{\prime},l^{\prime}=1}^{n}\sum_{i=1}^{d}\alpha_{i}\sum_{m=1}^{n}\big{(}(R_{n}(x_{i,j})^{*})^{q}\big{)}_{j^{\prime},m}\big{(}R_{n}(x_{i,l})^{q}\big{)}_{m,l^{\prime}}\mathrm{e}^{\mathrm{i}(j^{\prime}-l^{\prime})z}d_{l}(z)$
$=\sum_{i=1}^{d}\alpha_{i}\sum_{m=1}^{n}\bigg{|}\sum_{j=1}^{N}\sum_{j^{\prime}=1}^{n}d_{j}(z)\big{(}(R_{n}(x_{i,j}))^{q}\big{)}_{m,j^{\prime}}\mathrm{e}^{-\mathrm{i}j^{\prime}z}\bigg{|}^{2}\geq 0$
for $z\in\mathbb{T}$. In addition, let $\tilde{x}_{n,j}=S_{n}(R_{n}(x_{j}))$. Then, we have
$\bigg{(}\sum_{j,l=1}^{N}d_{j}^{*}k_{n}^{\operatorname{sep},q}(x_{j},x_{l})d_{l}\bigg{)}(z)$
$=\sum_{j,l=1}^{N}\overline{d_{j}(z)}\tilde{k}(\tilde{x}_{n,j},\tilde{x}_{n,l})d_{l}(z)\sum_{j^{\prime},l^{\prime}=1}^{n}\sum_{m_{1},\ldots,m_{2q-1}=1}^{n}$
$\big{(}(R_{n}(a)^{*})\big{)}_{j^{\prime},m_{1}}\big{(}(R_{n}(a)^{*})\big{)}_{m_{1},m_{2}}\cdots\big{(}(R_{n}(a)^{*})\big{)}_{m_{q-1},m_{q}}\big{(}R_{n}(a)\big{)}_{m_{q},m_{q+1}}\cdots\big{(}R_{n}(a)\big{)}_{m_{2q-1},l^{\prime}}\mathrm{e}^{\mathrm{i}(j^{\prime}-l^{\prime})z}$
$=\sum_{j,l=1}^{N}\overline{d_{j}(z)}\tilde{k}(\tilde{x}_{n,j},\tilde{x}_{n,l})d_{l}(z)$
$\qquad\sum_{m_{q}=1}^{n}\bigg{|}\sum_{j^{\prime}=1}^{n}\sum_{m_{1},\ldots,m_{q-1}=1}^{n}\big{(}R_{n}(a)\big{)}_{m_{q},m_{1}}\cdots\big{(}R_{n}(a)\big{)}_{m_{q-2},m_{q-1}}\big{(}R_{n}(a)\big{)}_{m_{q-1},j^{\prime}}\mathrm{e}^{-\mathrm{i}j^{\prime}z}\bigg{|}^{2}$
$\geq 0.$
Thus, $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{sep},q}$ are positive definite. $\Box$
Proposition 12    Let $\beta_{n}\geq-\min_{t\in\mathbb{T}^{q}}F_{n}^{q,P}(t)$. Let
$\hat{k}_{n}^{\operatorname{prod},q}(x,y)=k_{n}^{\operatorname{prod},q}(x,y)+\beta_{n}\int_{\mathbb{T}^{2q}}\prod_{j=1}^{q}\overline{\tilde{k}_{1,j}(x(t_{j}),y(t_{j}))}\tilde{k}_{2,j}(x(t_{q+j}),y(t_{q+j}))\mathrm{d}t.$
Then, $\hat{k}_{n}^{\operatorname{prod},q}$ is positive definite.
Proof  Let $x_{1},\ldots,x_{N}\in\mathcal{A}^{d}$ and $d_{1},\ldots,d_{N}\in\mathcal{A}$. Then, we have
$\bigg{(}\sum_{j,l=1}^{N}d_{j}^{*}k_{n}^{\operatorname{prod},q}(x_{j},x_{l})d_{l}\bigg{)}(z)$
$=\sum_{m,l=1}^{N}\int_{\mathbb{T}^{2q}}\prod_{j=1}^{q}\overline{d_{m}(z)}\overline{\tilde{k}_{1,j}(x_{m}(t_{j}),x_{l}(t_{j}))}\tilde{k}_{2,j}(x_{m}(t_{q+j}),x_{l}(t_{q+j}))d_{l}(z)F_{n}^{2q,P}(z\mathbf{1}-t)\mathrm{d}t$
$\qquad+\beta_{n}\sum_{m,l=1}^{N}\int_{\mathbb{T}^{2q}}\prod_{j=1}^{q}\overline{d_{m}(z)}\overline{\tilde{k}_{1,j}(x_{m}(t_{j}),x_{l}(t_{j}))}\tilde{k}_{2,j}(x_{m}(t_{q+j}),x_{l}(t_{q+j}))d_{l}(z)\mathrm{d}t$
$=\sum_{m,l=1}^{N}\int_{\mathbb{T}^{2q}}\prod_{j=1}^{q}\overline{d_{m}(z)}\overline{\tilde{k}_{1,j}(x_{m}(t_{j}),x_{l}(t_{j}))}\tilde{k}_{2,j}(x_{m}(t_{q+j}),x_{l}(t_{q+j}))d_{l}(z)(F_{n}^{2q,P}(z\mathbf{1}-t)+\beta_{n})\mathrm{d}t$
$\geq 0$
for $z\in\mathbb{T}$. The last inequality is derived since the map $([x_{1},x_{2}],[y_{1},y_{2}])\mapsto\overline{\tilde{k}_{1,j}(x_{1},y_{1})}\tilde{k}_{2,j}(x_{2},y_{2})$ is positive definite and $F_{n}^{2q,P}(z\mathbf{1}-t)+\beta_{n}\geq 0$. $\Box$
Lemma 13    We have $|F_{n}^{q,P}(z)|\leq n^{q}$.
Proof  The number of terms in $F_{n}^{q,P}$ is equal to the number of terms in $\sum_{m,l=0}^{n-1}(T_{1}\cdots T_{2q})_{m,l}$, where $T_{1},\ldots,T_{2q}$ are Toeplitz matrices. Thus, we have
$|F_{n}^{q,P}(z)|\leq\frac{1}{n}n^{q+1}=n^{q}.$
$\Box$
Proposition 16    Let $\mathcal{F}$ be an $\mathcal{A}$-valued function class. Assume for any $x\in\mathcal{A}_{0}^{d}$ and $z\in\mathbb{T}$, $f(x)(z)\in\mathbb{R}$. Then, we have $\hat{R}_{N}(\mathbf{x},\mathcal{F}(z))=\hat{R}_{\mathcal{A},N}(\mathbf{x},\mathcal{F})(z)$, where $\mathcal{F}(z)=\{x\mapsto(f(x))(z)\,\mid\,f\in\mathcal{F}\}$.
Proof  We first show $(\sup^{\mathcal{A}}_{a\in\mathcal{S}}a)(z)=\sup_{a\in\mathcal{S}}a(z)$ for $\mathcal{S}\subseteq\mathcal{A}_{+}$ for any $z\in\mathbb{T}$. Let $a\in\mathcal{S}$. Since $a(z)\leq\sup_{a\in\mathcal{S}}a(z)$ for any $z\in\mathbb{T}$, we have $a\leq_{\mathcal{A}}b$, where $b\in\mathcal{A}$ is defined as $b(z)=\sup_{a\in\mathcal{S}}a(z)$. Thus, $b$ is an upper bound of $\mathcal{S}$, and we have $\sup^{\mathcal{A}}_{a\in\mathcal{S}}a\leq_{\mathcal{A}}b$, which means $(\sup^{\mathcal{A}}_{a\in\mathcal{S}}a)(z)\leq\sup_{a\in\mathcal{S}}a(z)$ for any $z\in\mathbb{T}$. Conversely, since $a\leq_{\mathcal{A}}\sup^{\mathcal{A}}_{a\in\mathcal{S}}a$ for $a\in\mathcal{S}$, we have $a(z)\leq(\sup^{\mathcal{A}}_{a\in\mathcal{S}}a)(z)$ for any $z\in\mathbb{T}$. Thus, we have $\sup_{a\in\mathcal{S}}a(z)\leq(\sup^{\mathcal{A}}_{a\in\mathcal{S}}a)(z)$.
Therefore, we have
$\hat{R}_{\mathcal{A},N}(\mathbf{x},\mathcal{F})(z)$ $=\frac{1}{N}\mathrm{E}\bigg{[}\sup_{f\in\mathcal{F}}\!^{\mathcal{A}}\bigg{|}\sum_{i=1}^{N}f(x_{i})^{*}\sigma_{i}\bigg{|}_{\mathcal{A}}\bigg{]}(z)=\frac{1}{N}\mathrm{E}\bigg{[}\bigg{(}\sup_{f\in\mathcal{F}}\!^{\mathcal{A}}\bigg{|}\sum_{i=1}^{N}f(x_{i})^{*}\sigma_{i}\bigg{|}_{\mathcal{A}}\bigg{)}(z)\bigg{]}$
$=\frac{1}{N}\mathrm{E}\bigg{[}\sup_{f\in\mathcal{F}(z)}\bigg{|}\sum_{i=1}^{N}\overline{f(x_{i})}\sigma_{i}\bigg{|}\bigg{]}=\frac{1}{N}\mathrm{E}\bigg{[}\sup_{f\in\mathcal{F}(z)}\bigg{|}\sum_{i=1}^{N}f(x_{i})\sigma_{i}\bigg{|}\bigg{]}=\hat{R}_{N}(\mathbf{x},\mathcal{F}(z)),$
where the third equality is given by the identity $|a|_{\mathcal{A}}(z)=|a(z)|$, and the forth equality is satisfied since $\sigma_{i}$ is real-valued. $\Box$
Theorem 15    Assume $k^{\operatorname{poly},q}$, $k^{\operatorname{prod},q}$, and $k^{\operatorname{sep}q}$ are real-valued. Let $D(k_{n}^{\operatorname{poly},q},x)=\sum_{j=1}^{d}\alpha_{j}\|R_{n}(x_{j})\|^{2q}_{\operatorname{op}}$, $D(\hat{k}_{n}^{\operatorname{prod},q},x)=\prod_{j=1}^{q}(\|R_{n}(\tilde{k}_{1,j}(x,x))\|_{\operatorname{op}}\|R_{n}(\tilde{k}_{2,j}(x,x))\|_{\operatorname{op}})+\beta_{n}C$, and $D(k_{n}^{\operatorname{sep},q},x)=\tilde{k}(x,x)\prod_{j=1}^{q}\|R_{n}(a_{j})\|^{2}$ for $x\in\mathcal{A}_{0}^{d}$, where $\|\cdot\|_{\operatorname{op}}$ is the operator norm and $C=\prod_{j=1}^{q}\int_{\mathbb{T}}\tilde{k}_{1,j}(x(t),x(t))\mathrm{d}t\int_{\mathbb{T}}\tilde{k}_{2,j}(x(t),x(t))\mathrm{d}t$. Assume $\beta_{n}\leq\beta_{n+1}$ for $\hat{k}_{n}^{\operatorname{prod},q}$. For $k_{n}=k_{n}^{\operatorname{poly},q},\hat{k}_{n}^{\operatorname{prod},q},k_{n}^{\operatorname{sep},q}$ and for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
$\mathrm{E}[g(f(X),Y)]\leq_{\mathcal{A}}\frac{1}{N}\sum_{i=1}^{N}g(f(X_{i}),Y_{i})+2L\frac{B}{N}\bigg{(}\sum_{i=1}^{N}D(k_{n},X_{i})\bigg{)}^{1/2}1_{\mathcal{A}}+3\sqrt{\frac{\log 1/\delta}{N}}1_{\mathcal{A}}$
for any $f\in\mathcal{F}$. In addition, we have $D(k_{n},x)\leq D(k_{n+1},x)$.
According to Hashimoto et al. [14], the $\mathcal{A}$-valued Rademacher complexity for the RKHM associated with an $\mathcal{A}$-valued kernel $k$ is upperbounded by $\|k(x_{i},x_{i})\|_{\mathcal{A}}$ as follows.
Lemma 17
Let $B>0$ and =$\mathcal{F}=\{f\in\mathcal{M}_{k}\,\mid\,\|f\|_{k}\leq C\}$. Then, we have
$\hat{R}_{\mathcal{A},N}(\mathbf{x},\mathcal{F})\leq_{\mathcal{A}}\frac{B}{N}\bigg{(}\sum_{i=1}^{N}\|k(x_{i},x_{i})\|_{\mathcal{A}}\bigg{)}^{1/2}1_{\mathcal{A}}.$
We obtain the following lemma directly by Lemma 4.2 in [31].
Lemma 18
Let $g:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{+}$ be an error function. Assume there exists $L>0$ such that for $y\in\mathcal{A}_{1}$ and $z\in\mathbb{T}$, $x\mapsto g(x,y(z))$ is $L$-Lipschitz continuous. Assume also for any $x\in\mathcal{A}_{0}^{d}$ and $z\in\mathbb{T}$, $f(x)(z)\in\mathbb{R}$. Then, we have
$\frac{1}{N}\mathrm{E}\bigg{[}\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}g(f(x_{i})(z),y_{i}(z))\sigma_{i}\bigg{]}=\frac{1}{N}\mathrm{E}\bigg{[}\sup_{f\in\mathcal{F}(z)}\sum_{i=1}^{N}g(f(x_{i}),y_{i}(z))\sigma_{i}\bigg{]}\leq L\hat{R}_{N}(\mathbf{x},\mathcal{F}(z)).$
We apply the following lemma to obtain the inequality (3).
Lemma 19
Let $g$ be the same map defined in Lemma 18. For any $f\in\mathcal{F}$ and any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
$\mathrm{E}[g(f(X),Y)](z)\leq\frac{1}{N}\sum_{i=1}^{N}g(f(X_{i}),Y_{i})(z)+\frac{2}{N}\mathrm{E}_{\sigma_{i}}\bigg{[}\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}g(f(X_{i})(z),Y_{i}(z))\sigma_{i}\bigg{]}+3\sqrt{\frac{\log 1/\delta}{N}}.$
Proof  With probability at least $1-\delta$, we have
$\mathrm{E}[g(f(X),Y)](z)=\mathrm{E}[g(f(X)(z),Y(z))]$
$\quad\leq\frac{1}{N}\sum_{i=1}^{N}g(f(X_{i})(z),Y_{i}(z))+\frac{2}{N}\mathrm{E}_{\sigma_{i}}\bigg{[}\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}g(f(X_{i})(z),Y_{i}(z))\sigma_{i}\bigg{]}+3\sqrt{\frac{\log 1/\delta}{N}}$
$\quad=\frac{1}{N}\sum_{i=1}^{N}g(f(X_{i}),Y_{i})(z)+\frac{2}{N}\mathrm{E}_{\sigma_{i}}\bigg{[}\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}g(f(X_{i})(z),Y_{i}(z))\sigma_{i}\bigg{]}+3\sqrt{\frac{\log 1/\delta}{N}}.$
The inequality follows by Theorem 3.1 in [31]. $\Box$
The following lemma completes the proof of Theorem 15.
Lemma 20
Let $k_{n}$ be $k_{n}^{\operatorname{poly},q}$, $\hat{k}_{n}^{\operatorname{prod},q}$, or $k_{n}^{\operatorname{sep},q}$ defined in Section 3. Then, for $x\in\mathcal{A}^{d}$, $D(k_{n},x)\leq D(k_{n+1},x)$.
Proof  We evaluate $\|k_{n}(x,x)\|_{\mathcal{A}}$. For $k_{n}^{\operatorname{poly},q}$, we have
$\|k_{n}^{\operatorname{poly},q}(x,x)\|$ $=\bigg{\|}S_{n}\bigg{(}\sum_{i=1}^{d}\alpha_{i}(R_{n}(x_{i})^{*})^{q}R_{n}(y_{i})^{q}\bigg{)}\bigg{\|}_{\mathcal{A}}\leq\bigg{\|}\sum_{i=1}^{d}\alpha_{i}(R_{n}(x_{i})^{*})^{q}R_{n}(y_{i})^{q}\bigg{\|}_{\operatorname{op}}$
$\leq\sum_{i=1}^{d}\alpha_{i}\|R_{n}(x)\|^{2q}_{\operatorname{op}}.$
The first inequality is derived by the following inequalities for $A\in\mathbb{C}^{n\times n}$:
$|S_{n}(A)(z)|$ $=\bigg{|}\frac{1}{n}\sum_{j,l=0}^{n-1}A_{j,l}\mathrm{e}^{\mathrm{i}(j-l)z}\bigg{|}$
$=\bigg{|}\operatorname{tr}\bigg{(}\frac{1}{n}\mathbf{1}\operatorname{diag}(\mathrm{e}^{0},\ldots,\mathrm{e}^{\mathrm{i}(n-1)})A\operatorname{diag}(\mathrm{e}^{0z},\ldots,\mathrm{e}^{\mathrm{i}(n-1)z})\bigg{)}\bigg{|}$
$\leq\frac{1}{n}\|\mathbf{1}\|_{1}\|\operatorname{diag}(\mathrm{e}^{0},\ldots,\mathrm{e}^{\mathrm{i}(n-1)})A\operatorname{diag}(\mathrm{e}^{0z},\ldots,\mathrm{e}^{\mathrm{i}(n-1)z})\|_{\operatorname{op}}\leq\|A\|_{\operatorname{op}}$
for any $z\in\mathbb{T}$. Here, $\operatorname{tr}(A)$ is the trace of a matrix $A$, $\operatorname{diag}(a_{1},\ldots,a_{n})$ for $a_{1},\ldots,a_{n}\in\mathbb{C}$ is the diagonal matrix whose diagonal elements are $a_{1},\ldots,a_{n}$, $\mathbf{1}\in\mathbb{C}$ is the matrix whose elements are all $1$, and $\|A\|_{1}$ for $A\in\mathbb{C}^{n\times n}$ is the trace norm defined as $\|A\|_{1}=\operatorname{tr}((A^{*}A)^{1/2})$. We used the inequality $|\operatorname{tr}(AB)|\leq\|A\|_{1}\|B\|_{\operatorname{op}}$ for $A,B\in\mathbb{C}^{n\times n}$ [32, Chapter IX, Section 2].
In addition, there exists $v\in\mathbb{C}^{n\times n}$ such that $\|R_{n}(x)\|_{\operatorname{op}}=\|R_{n}(x)v\|$. Thus, we have
$\|R_{n}(x)\|_{\operatorname{op}}$ $=\|R_{n}(x)v\|=\|Q_{n}^{*}M_{x}Q_{n}v\|=\bigg{\|}Q_{n}^{*}M_{x}Q_{n+1}\begin{bmatrix}v\\ 0\end{bmatrix}\bigg{\|}\leq\bigg{\|}Q_{n+1}^{*}M_{x}Q_{n+1}\begin{bmatrix}v\\ 0\end{bmatrix}\bigg{\|}$
$\leq\|Q_{n+1}^{*}M_{x}Q_{n+1}\|_{\operatorname{op}}=\|R_{n+1}(x)\|_{\operatorname{op}},$
where $Q_{n}=[e_{1},\ldots,e_{n}]$ and $e_{j}(z)=\mathrm{e}^{\mathrm{i}jz}$ is the Fourier function. $\Box$
Appendix B Details of Remark 10
To understand the relationship between $q$ and the interactions, we observe the convergence shown in Lemma 9 from the perspective of $q$. The following lemma is by Brandlini and Travaglini [25].
Lemma 21
Let $H_{n}^{q,P}(t)=n^{q}\int_{0}^{1}\int_{\mathbb{R}^{q}}\chi_{jP}(r)\mathrm{e}^{\mathrm{i}nr\cdot t}\mathrm{d}r\mathrm{d}j$ for $t\in\mathbb{T}^{q}$, where $\chi_{S}$ is the characteristic function with respect to a set $S$. Then, we have $F_{n}^{q,P}(t)=\sum_{m\in\mathbb{Z}^{q}}H_{n}^{q,P}(t+2\pi m)$.
Proof  Let $u_{j,t}(r)=\chi_{jP}(r)\mathrm{e}^{\mathrm{i}r\cdot t}$. Then, we have
$F_{n}^{q,P}(t)$ $=\frac{1}{n}\sum_{j=1}^{n-1}\sum_{r\in jP\bigcap\mathbb{Z}^{q}}\mathrm{e}^{\mathrm{i}r\cdot t}=\frac{1}{n}\int_{0}^{n}\sum_{r\in\mathbb{Z}^{q}}\chi_{jP}(r)\mathrm{e}^{\mathrm{i}r\cdot t}\mathrm{d}j=\frac{1}{n}\int_{0}^{n}\sum_{m\in\mathbb{Z}^{q}}\widehat{u_{j,t}}(m)\mathrm{d}j$
$=\sum_{m\in\mathbb{Z}^{q}}\frac{1}{n}\int_{0}^{n}\int_{\mathbb{R}^{q}}\chi_{jP}(r)\mathrm{e}^{\mathrm{i}r\cdot t}\mathrm{e}^{2\pi\mathrm{i}r\cdot m}\mathrm{d}r\mathrm{d}j=\sum_{m\in\mathbb{Z}^{q}}\int_{0}^{1}\int_{\mathbb{R}^{q}}\chi_{jnP}(r)\mathrm{e}^{\mathrm{i}r\cdot t}\mathrm{e}^{2\pi\mathrm{i}r\cdot m}\mathrm{d}r\mathrm{d}j$
$=\sum_{m\in\mathbb{Z}^{q}}\int_{0}^{1}\int_{\mathbb{R}^{q}}\chi_{jP}\bigg{(}\frac{r}{n}\bigg{)}\mathrm{e}^{\mathrm{i}r\cdot t}\mathrm{e}^{2\pi\mathrm{i}r\cdot m}\mathrm{d}r\mathrm{d}j=n^{q}\sum_{m\in\mathbb{Z}^{q}}\int_{0}^{1}\int_{\mathbb{R}^{q}}\chi_{jP}(r)\mathrm{e}^{\mathrm{i}nr\cdot t}\mathrm{e}^{2\pi\mathrm{i}r\cdot m}\mathrm{d}r\mathrm{d}j,$
where $\hat{\cdot}$ is the Fourier transform, and the third equality is by the Poisson summation formula. $\Box$
To understand the convrgence, we split $F_{n}^{q,P}(t)=\sum_{m\in\mathbb{Z}^{q}}H_{n}^{q,P}(t+2\pi m)$ into two parts: $H_{n}^{q,P}(t)$ and $\tilde{H}_{n}^{q,P}(t):=\sum_{m\neq 0}H_{n}^{q,P}(t+2\pi m)$. In the convergence discussed in Lemma 9, the convolution $g\ast{H}_{n}^{q,P}(z)$ goes to $g(z)$. On the other hand, $g\ast\tilde{H}_{n}^{q,P}(z)$ goes to 00 [25, Theorem 1]. Regarding the convergence of $g\ast{H}_{n}^{q,P}(z)$ we can evaluate it as follows, especially for $q=2$.
Lemma 22
For $z\in\mathbb{T}$, we have
$\lim_{n\to\infty}g\ast H_{n}^{q,P}(z)-g(z)=\int_{0}^{1}\lim_{n\to\infty}\bigg{(}\int_{t\in\mathbb{T}^{q}}\int_{jnP}\mathrm{e}^{\mathrm{i}r\cdot(z-t)}g(t)\mathrm{d}r\mathrm{d}t-g(z)\bigg{)}\mathrm{d}j.$
Moreover, let $h^{P}_{s}(z)=\int_{t\in\mathbb{T}^{q}}\int_{sP}\mathrm{e}^{\mathrm{i}r\cdot(z-t)}g(t)\mathrm{d}r\mathrm{d}t-g(z)$. Let $q=2$, $Q=\{(r_{1},-r_{2})\,\mid\,(r_{1},r_{2})\in P\}$, $P_{1}^{\prime}=[-1,1]$, and $P_{2}=\{(r_{1},r_{2})\,\mid\,|r_{1}|\leq 1,\ |r_{2}|\leq 1,\ |r_{1}+r_{2}|\leq 1,\ |r_{1}-r_{2}|\leq 1\}$. Let $g(z)=\overline{x_{1}(z_{1})}\overline{x_{2}(z_{2})}y_{1}(z_{3})y_{4}(z_{4})$ for $z=[z_{1},z_{2},z_{3},z_{4}]$. Then, for $z\in\mathbb{T}^{2}$, there exists $C(z)\geq 0$ such that we have
$\frac{1}{2}|h^{P}_{jn}(z)+h^{Q}_{jn}(z)|\leq\frac{1}{2}C(z)\sum_{j=1}^{2}$ $\bigg{|}\int_{t\in\mathbb{T}}\int_{jnP_{1}^{\prime}}\mathrm{e}^{\mathrm{i}r(z_{1}-t)}x_{j}(t_{j})\mathrm{d}r\mathrm{d}t-x_{j}(z_{j})\bigg{|}$
$+\frac{1}{2}\bigg{|}\int_{\mathbb{T}^{2}}\int_{jnP_{2}}\mathrm{e}^{\mathrm{i}r\cdot(z-t)}g(t)\mathrm{d}r\mathrm{d}t_{1}-g(z)\bigg{|}.$ (6)
Proof  We have
$n^{q}\int_{t\in\mathbb{T}^{q}}\int_{0}^{1}\int_{jP}\mathrm{e}^{\mathrm{i}nr\cdot(z-t)}g(t)\mathrm{d}r\mathrm{d}j\mathrm{d}t-g(z)=\int_{0}^{1}\int_{t\in\mathbb{T}^{q}}\int_{jnP}\mathrm{e}^{\mathrm{i}r\cdot(z-t)}g(t)\mathrm{d}r\mathrm{d}t-g(z)\mathrm{d}j.$
Let ${h}^{P}_{s}(z)=\int_{t\in\mathbb{T}}\int_{sP}\mathrm{e}^{\mathrm{i}r\cdot(z-t)}g(t)\mathrm{d}r\mathrm{d}t-g(z)$. Then, $h^{P}_{0}(z)=0$, and since $\lim_{s\to\infty}h^{P}_{s}(z)=0$, there exists $D(z)>0$ such that $|h^{P}_{s}(z)|\leq D(z)$ for any $s>0$. Thus, $|h^{P}_{jn}(z)|\leq D(z)$ for any $n\in\mathbb{N}$ and $0<j\leq 1$. Thus, by the bounded convergence theorem, we have
$\lim_{n\to\infty}\int_{0}^{1}\int_{t\in\mathbb{T}^{q}}\int_{jnP}\mathrm{e}^{\mathrm{i}r\cdot(z-t)}g(t)\mathrm{d}r\mathrm{d}t-g(z)\mathrm{d}j=\int_{0}^{1}\lim_{n\to\infty}\int_{t\in\mathbb{T}^{q}}\int_{jnP}\mathrm{e}^{\mathrm{i}r\cdot(z-t)}g(t)\mathrm{d}r\mathrm{d}t-g(z)\mathrm{d}j.$
In addition, let $P_{1}=[-1,1]^{2}$ and $P_{2}=\{(r_{1},r_{2})\,\mid\,|r_{1}|\leq 1,\ |r_{2}|\leq 1,\ |r_{1}+r_{2}|\leq 1,\ |r_{1}-r_{2}|\leq 1\}$. Then, we have
$h^{P}_{jn}(z)+h^{Q}_{jn}(z)=h^{P_{1}}_{jn}(z)+h^{P_{2}}_{jn}(z).$
For $h^{P_{1}}_{jn}(z)$, we have
$|h^{P_{1}}_{jn}(z)|$ $=\bigg{|}\int_{\mathbb{T}}\int_{-jn}^{jn}\mathrm{e}^{\mathrm{i}r(z_{1}-t)}x_{1}(t)\mathrm{d}r\mathrm{d}t\int_{\mathbb{T}}\int_{-jn}^{jn}\mathrm{e}^{\mathrm{i}r(z_{2}-t)}x_{2}(t)\mathrm{d}r\mathrm{d}t-g(z)\bigg{|}$
$\leq\bigg{|}\int_{\mathbb{T}}\int_{-jn}^{jn}\mathrm{e}^{\mathrm{i}r(z_{1}-t)}x_{1}(t)\mathrm{d}r\mathrm{d}t\bigg{(}\int_{\mathbb{T}}\int_{-jn}^{jn}\mathrm{e}^{\mathrm{i}r(z_{2}-t)}x_{2}(t)\mathrm{d}r\mathrm{d}t-x_{2}(z_{2})\bigg{)}\bigg{|}$
$\quad+\bigg{|}x_{2}(z_{2})\bigg{(}\int_{\mathbb{T}}\int_{-jn}^{jn}\mathrm{e}^{\mathrm{i}r(z_{1}-t)}x_{1}(t)\mathrm{d}r\mathrm{d}t-x_{1}(z_{1})\bigg{)}\bigg{|}.$
In the same manner as $h_{s}^{P}(z)$, there exists $\tilde{C}(z)>0$ such that $|\int_{\mathbb{T}}\int_{-jn}^{jn}\mathrm{e}^{\mathrm{i}r(z_{1}-t)}x_{1}(t)\mathrm{d}r\mathrm{d}t|\leq\tilde{C}(z)$. By setting $C(z)=\max\{\tilde{C}(z),|x_{2}(z_{2})|\}$, we obtain the result. $\Box$
The left hand side is the average of $h_{jn}^{P}(z)$ and $h_{jn}^{Q}(z)$, where $P$ and $Q$ are symmetric with respect to the second coordinate $r_{2}$. The first term in the right hand side of Eq. (6) is described by the convergence of $h_{jn}^{[-1,1]}(z)$, which corresponds to the one-dimensional ($(q-1)$-dimensional) Fejér kernel. Since we have an additional term $1/2|\int_{\mathbb{T}^{2}}\int_{jnP_{2}}\mathrm{e}^{\mathrm{i}r\cdot(z-t)}g(t)\mathrm{d}r\mathrm{d}t-g(z)|$, if $C(z)>1$, the convergence of $h^{P}_{jn}(z)$ is expected to be slower than its one-dimensional counterpart $h^{[-1,1]}_{jn}(z)$. Based on this observation, we expect that the convergence in Theorem 7 becomes slower as $q$ becomes larger, and the interaction along $z\in\mathbb{T}$ becomes larger.
Appendix C Representer theorem for RKHMs over general $C^{*}$-algebras
Hashimoto et al. [33, Proposition 4.5] showed an approximate representer theorem for RKHMs over von Neumann-algebras. Since $C(\mathbb{T})$ is not a von Neumann-algebra, we generalize the theorem to that for general $C^{*}$-algebras as follows.
Proposition 23
Let $x_{1},\ldots,x_{N}\in\mathcal{A}^{d}$. Let $g:\mathcal{A}^{N}\to\mathcal{A}_{+}$ be a continuous map, let $h:\mathcal{A}_{+}\to\mathcal{A}_{+}$ satisfy $h(c)\leq_{\mathcal{A}}h(d)$ for $c,d\in\mathcal{A}_{+}$ with $c\leq_{\mathcal{A}}d$. Let $L(f)=g(f(x_{1}),\ldots,f(x_{N}))+h(|f|_{k})$ for $f\in\mathcal{M}_{k}$. If there exists a solution $f_{0}$ of the minimization problem $\min_{f\in\mathcal{M}_{k}}L(f)$, then for any $\epsilon>0$, there exists $\tilde{f}$ that admits the representation of $\tilde{f}=\sum_{i=1}^{N}\phi(x_{i})c_{i}$ and $\|L(\tilde{f})-L(f_{0})\|\leq\epsilon$.
Proof  For $f_{1},f_{2}\in\mathcal{M}_{k}$, let $\theta_{f_{1},f_{2}}:\mathcal{M}_{k}\to\mathcal{M}_{k}$ defined as $\theta_{f_{1},f_{2}}(v)=f_{1}\left\langle f_{2},v\right\rangle_{k}$ for $v\in\mathcal{M}_{k}$. Let $\mathcal{N}_{k}$ be the submodule generated algebraically by $\{\phi(x_{1}),\ldots,\phi(x_{N})\}$, and let $\mathcal{B}=\overline{\operatorname{Span}\{\theta_{f_{1},f_{2}}\,\mid\,f_{1},f_{2}\in\mathcal{N}_{k}\}}$. In addition, let $\mathcal{L}(\mathcal{M}_{k})$ be the (unital) $C^{*}$-algebra of adjointable $\mathcal{A}$-linear operators on $\mathcal{M}_{k}$ (see [12, p8]). Since $\mathcal{B}$ is a $C^{*}$-subalgebra of $\mathcal{L}(\mathcal{M}_{k})$ (Note $\theta_{f_{1},f_{2}}\theta_{f_{3},f_{4}}=\theta_{f_{1},f_{4}\left\langle f_{3},f_{2}\right\rangle}$ and $\theta_{f_{1},f_{2}}^{*}=\theta_{f_{2},f_{1}}$), it has a net $\{b_{i}\}_{i}$ such that $10\leq_{\mathcal{L}(\mathcal{M}_{k})}b_{i}\leq_{\mathcal{L}(\mathcal{M}_{k})}$ 1 and $\lim_{i}\theta_{f_{1},f_{2}}b_{i}=\lim_{i}b_{i}\theta_{f_{1},f_{2}}=\theta_{f_{1},f_{2}}$ for any $f_{1},f_{2}\in\mathcal{N}_{k}$ [23, Theorem I.4.8]. For each $v\in\mathcal{N}_{k}$, we have
$\|v-b_{i}v\|_{k}^{2}$ $=\|\left\langle(1-b_{i})v,(1-b_{i})v\right\rangle_{k}\|_{\mathcal{A}}=\|\theta_{(1-b_{i})v,(1-b_{i})v}\|_{\mathcal{L}(\mathcal{M}_{k})}$
$=\|(1-b_{i})\theta_{v,v}(1-b_{i})^{*}\|_{\mathcal{L}(\mathcal{M}_{k})}=\|(1-b_{i})\theta_{v,v}(1-b_{i})\|_{\mathcal{L}(\mathcal{M}_{k})}.$
Thus, we have $\lim_{i}b_{i}v=v$. Since $g$ is continuous, we have
$\lim_{i}g(\left\langle\phi_{k}(x_{1}),b_{i}f_{0}\right\rangle_{k},\ldots,\left\langle\phi_{k}(x_{N}),b_{i}f_{0}\right\rangle_{k})=\lim_{i}g(\left\langle b_{i}^{*}\phi_{k}(x_{1}),f_{0}\right\rangle_{k},\ldots,\left\langle b_{i}^{*}\phi_{k}(x_{N}),f_{0}\right\rangle_{k})$
$=\lim_{i}g(\left\langle b_{i}\phi_{k}(x_{1}),f_{0}\right\rangle_{k},\ldots,\left\langle b_{i}\phi_{k}(x_{N}),f_{0}\right\rangle_{k})=g(\left\langle\phi_{k}(x_{1}),f_{0}\right\rangle_{k},\ldots,\left\langle\phi_{k}(x_{N}),f_{0}\right\rangle_{k}).$
Therefore, for $\epsilon>0$, there exists $i$ such that
$\|g((b_{i}f_{0})(x_{1}),\ldots,(b_{i}f_{0})(x_{N}))-g(f_{0}(x_{1}),\ldots,f_{0}(x_{N}))\|_{\mathcal{A}}\leq\epsilon.$
Since $b_{i}\geq_{\mathcal{L}(\mathcal{M}_{k})}0$, there exists $c\geq_{\mathcal{L}(\mathcal{M}_{k})}0$ such that $b_{i}=c^{2}$. Thus, $b_{i}-b_{i}^{2}=c^{2}-c^{4}=c(1-c^{2})c=c^{*}(1-b_{i})c\geq_{\mathcal{L}(\mathcal{M}_{k})}0$. Thus, $10\leq_{\mathcal{L}(\mathcal{M}_{k})}b_{i}^{2}\leq_{\mathcal{L}(\mathcal{M}_{k})}b_{i}\leq_{\mathcal{L}(\mathcal{M}_{k})}$ 1, and we have $|b_{i}f_{0}|_{k}^{2}=\left\langle f_{0},b_{i}^{2}f_{0}\right\rangle_{k}\leq_{\mathcal{A}}\left\langle f_{0},f_{0}\right\rangle_{k}=|f_{0}|_{k}$. As a result, we obtain
00 0 $\leq_{\mathcal{A}}L(b_{i}f_{0})-L(f_{0})$
$=g((b_{i}f_{0})(x_{1}),\ldots,(b_{i}f_{0})(x_{N}))+h(|b_{i}f_{0}|_{k})-g(f_{0}(x_{1}),\ldots,f_{0}(x_{N}))-h(|f_{0}|_{k})$
$\leq_{\mathcal{A}}g((b_{i}f_{0})(x_{1}),\ldots,(b_{i}f_{0})(x_{N}))-g(f_{0}(x_{1}),\ldots,f_{0}(x_{N})).$
Since $b_{i}f_{0}\in\mathcal{N}_{k}$, setting $\tilde{f}=b_{i}f_{0}$ completes the proof of the proposition. $\Box$
Appendix D Experimental Details and Additional Results
We provide experimental details below. All the experiments in this paper were executed with Python 3.9 on an Intel(R) Core(TM) i9-10885H 2.4GHz processor with the Windows 10 operating system.
Figure 3: Eigenvalues of the Gram matrix $\mathbf{G}(0)$ indexed as the descending order. (Average value of results of five different runs. The error bar represents the standard deviation.)
D.1 Experiment with synthetic data
We estimated $f$ using kernel ridge regression with the regularization parameter $\lambda=0.01$. We generated 1000 input test samples in the same manner as the input training samples, and evaluated the test error $1/N\sum_{i=1}^{N}\|\hat{f}(\tilde{x}^{i})-\tilde{y}^{i}\|_{L^{2}(\mathbb{T})}$, where $\hat{f}$ is the estimation of $f$ and $\tilde{x}^{i}$ and $\tilde{y}^{i}$ are input and output test samples, respectively. For $x^{i}$, $y^{i}$, $\tilde{x}^{i}$, and $a$, we discretized each function with 30 equally spaced points on $\mathbb{T}$. In addition, to investigate the positive definiteness of $\hat{k}_{n}^{\operatorname{prod},q}$ discussed in Remark 14, we computed the eigenvalues of the Gram matrix $\mathbf{G}(0)$ at 0. Figure 3 shows the result. The index $i$ is determined in the descendent order with respect to the value of $\lambda_{i}$. Thus, we can see that although we set $\beta_{n}$ as a small value, all the eigenvalues of $\mathbf{G}(0)$ are positive in this case.
D.2 Experiment with MNIST
We generated 200 input test samples in the same manner as the input training samples, and evaluated the test error $1/N\sum_{i=1}^{N}\|\hat{f}(\tilde{x}^{i})-\tilde{y}^{i}\|_{L^{2}(\mathbb{T})}$, where $\hat{f}$ is the estimated map that maps the image with the missing part to its original image, and $\tilde{x}^{i}$ and $\tilde{y}^{i}$ are input and output test samples, respectively.